JT:

Person 1 can do a certain job in twenty-one minutes, and person 2 can do the same job in twenty-eight minutes. When completing the job together, which expression would be used to represent the amount of work done by person 1? 21 1/21 x/21 21x

Hero:

In general, when trying to solve work problems, what I usually do is begin with this formula: \(\dfrac{AB}{A + B} = t\) Where A represents time it takes the 1st person to do the job alone B represents time it takes the 2nd person to do the job alone t represents the time it takes both to do the job together. However in this case, they want you to set it up this way: \(\dfrac{1}{A} + \dfrac{1}{B} = \dfrac{1}{t}\)

Hero:

@JT

JT:

I think that it's x/21

JT:

@Hero

Hero:

Not correct